OpenStudy (anonymous):

Can anyone give me practice problems for Simplifying Numerical Expressions with Rational Exponents?

OpenStudy (dumbcow):

hold on

OpenStudy (anonymous):

\[32^{\frac{ 1 }{ 5 }}\] \[81^{\frac{ 3 }{ 4 }}\] \[125^{\frac{ -2 }{ 3 }}\]

OpenStudy (anonymous):

Thank you!